hello please help me with the following equation:
22-5(6v-1)=-63
I have gotten the answer of v=3 from other people on the internet but i dont have a clue how they did it and it makes no sense to me so please explain thank youu!

Answer by rfer(16322)

22-5(6v-1)=-63
22-30v+5=-63
30v=-63-27
30v=90
v=3